Katalin and Patrick both decide to visit their friend Moti who lives 90 miles away. Patrick drives at a rate of 45 mi/h and Katalin drives at a rate of 55 mi/h. How far will Patrick still have to travel by the time Katalin reaches Moti's home?
Answer by checkley71(8403)

90/55=1.64 hours for Katalin.
45*1.64=73.8 miles traveled by Patrick in 1.64 hours.
90-73.8=16.2 miles left for Patrick when Katalan arrived at Moti's.